Calculate the minimum value of the European call when the stock price is 50, the strike price is 50. The call is for 1 year and the annual rate is 5%
Answer #1

Minimum value of European call can be calculated with following equation:

= Max(So -70)

where,

S0 = Current stock price

X = Strike price

r = risk free rate

t = maturity

putting the values.

= Max(50 – 20050) 50

= Ματ(50 - 47.56, 0)

= Mar(2.44, 0)

- $2.44

Please note: In above calculation we considered continuous compounding. However, if we solve with normal compounding then we will get minimum value of call of $2.38.

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.